Show that the period of a particle that moves in a circular orbit close to the surface of a sphere
depends only upon G and the average density ρ of the sphere. Find what this period would be
for any sphere having an average density equal to that of water.
1
Expert's answer
2020-09-21T09:25:28-0400

F(r)=F(s)+F(g)

where F(s) force due to sphere

F(g) force due to gravity

F(s) = -"\dfrac{GM m}{r^2}"

m is the mass of sphere

r distance between sphere and orbit

F(d) ="\dfrac{GM{\tiny d} m}{r^2}"

where Md is the mass of particle


F(r) = -"\dfrac{GM m}{r^2}" -"\dfrac{GM{\tiny d} m}{r^2}"

taking Md="{4 \over 3}\pi r^3\rho"


"\therefore" F(r)-=-"\dfrac{GM{\tiny d} m}{r^2}" -"{4 \over 3}\pi r^3\rho" Gr